Rationalise the denominator 1/2+√3.

Grade:9

2 Answers

vamsi vamsi
9 years ago
1(2-√3)/(2+√3)(2-√3)=(2-√3)/4-(√3)2=2-√3/1.
Arun Kumar IIT Delhi
9 years ago
Hello

Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty